### v2

# Accepted with Minor Revisions: Value of AI-Assisted Scientific Writing

Abstract.
Large Language Models have seen expanding application across domains, yet their effectiveness as assistive tools for scientific writing—an endeavor requiring precision, multimodal synthesis, and domain expertise—remains insufficiently understood. We examine the potential of LLMs to support domain experts in scientific writing, with a focus on abstract composition. We design an incentivized randomized controlled trial with a hypothetical conference setup where participants with relevant expertise are split into an author and reviewer pool. Inspired by methods in behavioral science, our novel incentive structure encourages authors to edit the provided abstracts to an acceptable quality for a peer-reviewed submission. Our 2 x 2 between-subject design expands into two dimensions: the implicit source of the provided abstract and the disclosure of it. We find authors make most edits when editing human-written abstracts compared to AI-generated abstracts without source attribution, often guided by higher perceived readability in AI generation. Upon disclosure of source information, the volume of edits converges in both source treatments. Reviewer decisions remain unaffected by the source of the abstract, but bear a significant correlation with the number of edits made. Careful stylistic edits, especially in the case of AI-generated abstracts, in the presence of source information, improve the chance of acceptance. We find that AI-generated abstracts hold potential to reach comparable levels of acceptability to human-written ones with minimal revision, and that perceptions of AI authorship, rather than objective quality, drive much of the observed editing behavior. Our findings reverberate the significance of source disclosure in collaborative scientific writing.
Generative AI, Human-AI Writing Assistance, Large Language Models, Scientific Writing, Text Editing and Evaluation, Alignment, Behavioral Science, Incentivized Random Trial
journalyear: 2026copyright: ccconference: 31st International Conference on Intelligent User Interfaces; March 23–26, 2026; Paphos, Cyprusbooktitle: 31st International Conference on Intelligent User Interfaces (IUI ’26), March 23–26, 2026, Paphos, Cyprusdoi: 10.1145/3742413.3789140isbn: 979-8-4007-1984-4/2026/03ccs: Human-centered computing Empirical studies in HCI
1. Introduction
Telling people about research is as important as doing it. Scientific publishing is an integral part of science, meant to disseminate research findings, foster collaboration, encourage reproducibility, and ensure that scientific knowledge is accessible and built upon over time (Liang et al., 2024). Bad writing can and often does prevent or delay the publication of good science (Gastel and Day, 2022). Researchers and scientists today, however, are often not formally trained to write, and generally learn by imitating the styles of their advisors and other authors (Lindsay, 2020). Large language models (LLMs), with their writing abilities, offer a promising venue to assist researchers in both conducting and writing about research. However, prior work on thorough assessment of their capabilities in open ended scientific writing tasks remains limited.
Figure 1. Curious case of scientific writing: We find that expert authors (e.g., with PhD) make most edits when provided with human-written counterparts of AI-generated abstracts, especially so when the source of abstracts remains unattributed. With attribution, we see an opposite trend: authors made careful stylistic edits when the abstract was known to be AI-generated, which often raises the chance of getting the abstract accepted.
Even so, the use of language models continues to grow in this area. Many tools and platforms termed “AI co-scientist” have emerged in the past year (Gottweis et al., 2025; Beel et al., 2025). They are marketed to be capable of generating, executing, and writing papers, with some of them getting accepted in academic workshops organized by peer-reviewed conferences (Yamada et al., 2025). Popular scientific writing platforms such as Overleaf now offer writing assistant plugins. [^1] https://www.overleaf.com/learn/how-to/Writefull_integration However, anecdotal examples of both seemingly AI-generated papers (Cybernews, 2023; DeGeurin, 2024) and peer reviews (Oransky and Marcus, 2024) have inspired criticisms (Liang et al., 2024). In addition, across disciplines, recent studies have raised concerns over the declining quality of writing in scientific publications in several disciplines, attributing it to the rise of LLM-generated content (Matsui, 2024; Lemire, 2024; Geng and Trotta, 2024). These studies, however, have largely focused on corpus-level trends on unreviewed manuscripts and are conducted without consulting or involving writers such as the authors of said publications.
To raise awareness and prevent misuse by bad actors or unintentional use by over-reliant writers, there is an urgent need to conduct a systematic evaluation of the writing abilities of the current state-of-the-art models in specialized domains. Furthermore, despite current limitations, AI assistance holds tremendous potential in accelerating future scientific progress (Sourati and Evans, 2023; Nathani et al., 2025; Majumder et al., 2024a). Towards understanding scientific writing capabilities, we seek to answer the following research question in this work: maintaining the standard for scientific writing quality, to what extent do researchers rely on model-generated content for academic publishing?
To answer this question, we simulate a realistic writing environment where authors (domain experts), the intended users of an AI assistant, use and edit model-generated text to create scientific prose with the goal of getting it accepted by reviewers (also domain experts) for publication. To scope this work, we focus on the task of writing abstracts. An abstract is meant to summarize research findings, highlight key results, and pique the readers’ interest. Given selected research findings, we evaluate how existing state-of-the-art LLMs perform at generating abstract drafts.
We simulate an abstract editing and reviewing process mimicking the process of scientific publication in a laboratory setting using the methodology of incentivized behavioral experiments (Azrieli et al., 2018) to encourage participants to perform the task to the best of their ability. Authors edit the abstracts, and we quantify author reliance based on the extent of their modifications to make them submission-ready for a hypothetical conference publication. The task is conducted under two source information conditions: (1) authors are informed that the original abstract provided to them was AI-generated, and (2) authors are not informed that the abstract was assembled by generative AI. To contextualize this behavior with a more realistic workflow, we also provide the authors with human-written abstracts under the same source information conditions. By considering abstracts from different sources (human experts and LLMs), we assess the reception of LLM-generated abstracts by evaluating correlations between acceptance rates, edit frequency, and edit types.
The edited abstracts are then evaluated by a group of (incentivized) reviewers in a double-blind fashion, who vote to accept or reject the submitted abstract. This design allows us to examine how authors’ reliance varies across source types and source information conditions, and how expert reviewers perceive the quality and credibility of abstracts produced through generative AI workflows. These experiments result in a collection of model-generated and human-edited texts, where we capture keystroke-level edits. From this rich data of inputs, abstracts, and edits, we conduct several evaluations and analyses. Our main evaluation is based on the correlation between the rate of editing and the final rating an abstract receives, contrasting the controls and different treatments. We consider and measure the impact of variables like writer expertise (undergraduate students, graduate students, experienced researchers), and demographics and social factors (gender, perception of AI). In addition, we quantitatively (using linguistic tools) and qualitatively (via interviews with participants) characterize different kinds of edits authors make to understand the differences between edits made to AI- and human-authored abstracts.
Across our experiments, we find that both the implicit origin of abstract and the disclosure of it systematically shape editing behavior and outcomes. Without source disclosure, authors make fewer edits to AI-generated abstracts, perceiving them as more readable, though PhD-level authors edited AI text more extensively. Despite these differences, reviewer evaluations showed no differences in accepting abstracts edited under no disclosure conditions. Disclosure of source primarily triggers social adjustments: authors reduced edits for human-written text when the source was revealed, whereas disclosure had only minor effects on AI-generated abstracts, with small impacts on perceived readability and confidence.
Our quantitative analysis of stylistic metrics reveals that edits to AI-generated abstracts are to improve cohesion, reduce nominalizations, and produce informative sentences. On the contrary, authors tend to lengthen the openings of the human-written abstracts. Similarly, disclosure of source alone has a minimal direct effect on the writing style. Finally, thematic analysis of our interviews with authors confirms that authors adopt different strategies depending on the origin of the abstracts: they simplify AI-generated text while restructuring human-authored abstracts for emphasis. Disclosure promotes a sense of accountability for both human- and AI-authored texts. Together, through a comprehensive large-scale experiment, we first highlight the varying cognitive editorial strategies emerging from writing that involves both human and AI-generated scientific writing. [^2] Our experimental data, task instructions, and example of a task are provided as supplementary.
2. Related Works
2.1. AI Systems for Automating Scientific Research
Early works on AI-assistance in scientific research largely focused on surface-level corrections and sentence-level improvements in writing (Ito et al., 2019; Bezerra et al., 2021) with some attempts at drafting entire paper structure (Wang et al., 2019). Recent work has started to apply LLMs to the entire research cycle, including hypothesis discovery (Zhou et al., 2024; Zhang et al., 2025; Majumder et al., 2024b; Abdel-Rehim et al., 2025), experiment planning and execution (Li et al., 2025; Binbas, 2025; Gui and Kim, 2025), scientific writing (Kanna, 2024; Liebling et al., 2025; Kousha and Thelwall, 2025), and writing reviews (Silva et al., 2025). Within the writing stage, some works have focused on specific subtasks like metaphor generation (Kim et al., 2023), figure caption generation (Kim et al., 2025), citation generation (Xing et al., 2020; Fierro et al., 2024; Jung et al., 2022), and related work generation (Funkquist et al., 2022; Li and Ouyang, 2024). Such tasks are, typically, accompanied by reference answers, making evaluation relatively straightforward. More recent work has started focusing on generating entire papers without any human intervention (Lu et al., 2024). In this work, we adopt a more realistic setup of humans relying on an AI-assistant to generate drafts, which they then edit (Kanna, 2024). We also use a principled method of human evaluation in the form peer review conducted via incentivized experiments.
A wide range of commercial [^3] https://www.gatsbi.com/, https://anara.com/blog/ai-research-tools as well as open-source tools now offer services for every stage of the writing process, from literature review and synthesis to full manuscript drafting, editing, and refinement (Shao et al., 2025). Specific features, such as automated abstract and title generators, are also becoming commonplace. [^4] https://www.writefull.com/ The widespread availability and marketing of these tools signify a de facto adoption that necessitates rigorous empirical evaluation of their effectiveness, their impact on the quality of scientific output, and the human behaviors they engender, which our study addresses. In fact, large-scale corpus analyses suggest that AI assistance in scientific writing is already widespread (Kobak et al., 2025) in certain domains. Recent studies emphasize the need to understand how human researchers interact with AI-generated drafts, how much agency they retain, and how these systems reshape scientific communication norms (Reza et al., 2025). Our work contributes to this line of inquiry by focusing on human reliance and editorial behavior when interacting with AI-generated scientific text under realistic publication incentives.
2.2. Human AI Collaboration and Co-writing
The majority of prior research on human-AI co-writing has focused on creative writing (Singh et al., 2025). Using LLMs, researchers have developed and evaluated systems supporting story writing (Yuan et al., 2022), playwriting (Grigis and De Angeli, 2024), and character development (Park et al., 2025). Some works also explore higher-level writing tasks such as prewriting (Wan et al., 2024), and generating perspective-specific feedback (Rashkin et al., 2025). Specialized creative applications have emerged for tasks including metaphor generation (van Heerden and Bas, 2024), collaborative storytelling (Wang et al., 2024), and personal diary writing (Kim et al., 2024). User studies reveal complex dynamics in how writers interact with and perceive AI writing assistance. At a system interaction level, the design of AI suggestions significantly impacts user behavior and output: sentence-level suggestions promote original content creation, while paragraph-level suggestions improve efficiency (Fu et al., 2023; Dhillon et al., 2024). Writers’ engagement with AI assistance is also influenced by their personal values and goals (Biermann et al., 2022; Guo et al., 2025). Writers show varying receptivity to AI support based on their confidence levels, demonstrating higher acceptance in areas where they lack expertise (Al-Bukhrani et al., 2025; Wang and Lu, 2025), and their desires for support are closely tied to their perception of support actors and personal values (Hwang et al., 2025; Guo et al., 2024). Professional writers note persistent challenges with AI systems’ ability to maintain consistent style and voice (Chakrabarty et al., 2024a).
Moreover, this human-AI writing relationship raises important concerns. Studies reveal that biased AI models can influence not only the resulting text but also users’ own opinions (Jakesch et al., 2023). Prior research has identified biases, quality issues, lack of structure, and superficiality in AI-generated text, especially in domains like fiction and satire (Chakrabarty et al., 2024a, b; Ippolito et al., 2022; Marco et al., 2024; Mirowski et al., 2023, 2024). Researchers have also found LLM-generated content to be overly positive and lacking in nuance and have highlighted the need for more diverse and representative training data to mitigate these shortcomings (Tian et al., 2024). Our work adapts evaluation practices from prior studies in co-writing by focusing on the quality of AI-generated text using the amount of edits made to it by a human co-writer. Related to our work, experimental studies demonstrate that higher-level AI scaffolding can improve writing quality and efficiency, particularly for less experienced writers, while simultaneously raising concerns around diminished authorship and ownership (Dhillon et al., 2024). In parallel, large-scale reviews of human–AI co-writing systems emphasize that writers desire different forms of AI support at different stages of the writing process, such as ideation, drafting, and revision (Reza et al., 2025). This underscores the importance of evaluating AI assistance within realistic workflows rather than isolated generation tasks.
2.3. Cognitive Biases in Human-AI interaction
The inability of humans to reliably detect AI-generated text makes their beliefs about a text’s origin a powerful psychological variable, opening the door for cognitive biases to influence their judgment and behavior (Jakesch et al., 2022; Henestrosa and Kimmerle, 2024; Fiedler and Döpke, 2025). A rich literature in HCI and psychology documents two opposing, context-dependent human tendencies when interacting with automated systems (Parasuraman et al., 2000; Lee and See, 2004; Dzindolet et al., 2003). On one hand, automation bias describes the tendency to over-rely on or excessively trust automated outputs, often using them as a cognitive shortcut to reduce mental effort (Mosier and Skitka, 2018; Lyell and Coiera, 2017). This can lead to errors of commission (accepting incorrect AI-generated information) and errors of omission (failing to notice problems that the AI missed) (Abdelwanis et al., 2024).
On the other hand, algorithm aversion describes a tendency to prefer human judgment over algorithmic judgment, even in cases where the algorithm is demonstrably better or equivalent (Sunstein and Gaffe, 2025). This aversion is often heightened in high-stakes, subjective, or ethically charged tasks, such as medical diagnosis (Filiz et al., 2023). It can be driven by a fundamental desire for human agency, a negative emotional reaction to being judged by a machine, and a belief that human experts possess unique, ineffable knowledge that algorithms cannot capture.
Our study advances the literature by shifting the focus from perceptual measures of AI-generated text such as self-reported “credibility” or “trust” (Lermann Henestrosa et al., 2023; Graefe and Bohlken, 2020) to concrete, measurable behavior in terms of amount of edits made to the text. The finding that attribution significantly impacts the number of edits demonstrates that this cognitive bias translates into tangible, effortful action on the part of the expert, also mirrored by recent empirical work (Zhu et al., 2025). Controlled experiments in German theses by Fiedler and Döpke (2025) show that when identical texts are labeled as AI-generated versus human-written, evaluators systematically favor the human-labeled versions, even when the labels are incorrect, demonstrating a strong source-identity bias against AI authorship. Conversely, when source information is withheld, both expert and non-expert evaluators struggle to reliably distinguish AI-generated academic writing from human writing, often rating them as comparable in quality.
2.4. Peer Review and Scholarly Evaluation
Our experimental design is situated within the broader field of science of peer review, which treats scholarly evaluation not as a purely objective process, but as a human endeavor subject to cognitive and social biases. A significant body of literature has used experimental methods to uncover these biases. For instance, landmark studies have shown that reviewer recommendations are significantly influenced by the prestige of an author’s institution, a bias that double-blinding is intended to mitigate (Ceci and Peters, 1982; Tomkins et al., 2017). Relatedly, we study the perceived origin of the text itself as a bias. The attribution of an abstract to either a human or an AI serves as an analogue to author identity, allowing us to investigate whether a “source identity bias” influences the authorial and review process in a similar manner to established forms of author bias.
In this work, following the methodology of behavioral scientists, we conduct incentivized randomized controlled trials of increasing complexity for this evaluation. Randomized controlled trials are widely regarded as the gold standard for establishing causal relationships in behavioral interventions (Banerjee and Duflo, 2009). Monetary incentives promote task engagement and data reliability. Participants align the study’s rewards with their decisions through a cost-benefit analysis, yielding behavior that closely mirrors real-world decision-making (Gneezy and Rustichini, 2000). Recent studies have employed RCTs to assess the effectiveness of AI tools in various domains, including writing and education (Kizilcec and Halawa, 2015). These trials provide empirical evidence on how AI interventions influence user outcomes, supporting their utility in scientific research.
Figure 2. Experimental workflow: Authors from a shared pool were randomly assigned to one of four treatments varying abstract source (AI-generated vs. human-written) and information disclosure (with vs. without source information). Authors revised abstracts in the editing interface. Each edited abstract was randomly assigned to three reviewers for their individual verdict. Reviewers only see the final edited version of the abstracts. A majority vote decides the final verdict.
3. Experiment Design
Writing scientific prose is a challenging task. It is intended to communicate complex ideas to the academic community or to the public, focusing on preciseness, clarity, and brevity (Lindsay, 2020). It requires the synthesis of diverse specialized knowledge and information sources, including non-textual elements like tables, statistics, and figures, alongside long contexts with strict formatting requirements. Performed by domain experts such as researchers with varying expertise, it is an inherently iterative process (Mao et al., 2019; Cornell, 2021). It is often collaborative, involving people with different experiences contributing, editing, and commenting on each other’s written work.
We aim to create a test-bed to evaluate how humans perceive and evaluate scientific writing. We follow a 4-step approach: (1) We select original research publications from peer-reviewed conference publication venues and curate AI-generated abstracts from the content of these papers using state-of-the-art LLMs; (2) We engineer an incentivized hypothetical conference set-up where we recruit human participants online to form an Author and a Reviewer pool. This allows us to emulate a real-world writing setting for both parties to play their part in the study; (3) We develop a web platform based on FirePad [^5] The editing platform codebase is available at: https://github.com/skai-research/scientific-writing-assistance to capture fine-grained character-level edits such as insertion, deletion, or substitution with timestamps. Each abstract is shown to the authors on this platform to generate edited abstracts. (4) We show each edited abstract to three independent reviewers to obtain the final accept/reject decision by majority voting.
3.1. Collecting and Curating Abstracts
For the scope of this study, we focus on the computer science (CS) domain for several reasons. First, CS is a fast-paced field with high publication volume (CVPR 2024, 2024), rapid turnaround cycles, and a high impact factor, particularly in conferences (Reuters, 2024; Freyne et al., 2010). This makes it a relevant and timely setting to study the integration of AI in scientific writing. Second, the widespread adoption of AI tools by researchers in CS provides ecological validity for examining human-AI collaboration in this context. Finally, as CS researchers ourselves, we are well-positioned to assess the accuracy and relevance of domain-specific content. While our study centers on CS, the design is generalizable and can be extended to other scientific domains. A typical research process involves identifying a research problem, reviewing existing literature, formulating a hypothesis, collecting data, analyzing the data, and potentially modifying the hypothesis and repeating the process. At the end, the researchers draw conclusions and present the findings. We evaluate the role of AI-assistance in the final stage, where, given research findings, an LLM generates a coherent abstract summarizing the results of the study in a factual, coherent, and appealing way.
Table 1. Abstracts from papers across venues
Venue Count Average Citation Year
NeurIPS 6 18.83 2024
CVPR 6 24.17 2024
ACL 6 30.33 2024
SIGCOMM 6 5.83 2024
VLDB 5 2.60 2024
IEEE S&P 6 19.83 2024
ICSE 5 10.80 2024
ASPLOS 5 15.40 2024
The most realistic setting to conduct our experiment is to involve different researchers who are at the end of their studies and wish to write up their findings. This setting, however, is expensive and prohibitively slow to emulate. Instead, we choose already published recent papers and convert them into a format of “research excerpts” that could be fed to an LLM to generate abstracts. We deliberately choose already published papers to isolate the effect of the quality of writing from the quality of the research itself; we only aim to evaluate the latter.
We select 45 published papers from top-tier conference venues in CS. [^6] To ensure that our selected research papers were not a part of the training dataset of the state-of-the-art LLMs, we use a subset of only very recent papers published in 2024 in the top CS conferences. For details on selected papers, please refer to Table 3. We mainly focus on application-based papers from subfields of CS, such as NLP, HCI, Systems, and others. [^7] We refrain from using lengthy papers with theoretical underpinnings, as their heavy dependence on background knowledge may make it challenging to generate concise research excerpts without losing information. For each venue’s latest publication, we randomly sample papers as shown in LABEL:tab:venuelist. Using GPT-4o, we create research excerpts, i.e., extract all relevant text from the paper that is necessary to write an abstract. [^8] OpenAI’s GPT-4o stood as the most recent model available during the time of the experiment. Our conclusions are conditional on this temporal choice and do not presume continued popularity or dominance of the model or its provider. We provide the exact prompt we use in Figure 13 (Appendix). Note, we focus on extraction rather than generation to make sure no facts in the original paper are misrepresented in the generated plan. We also manually verify each model output for factual correctness. We provide a sample output in Figure 14 (Appendix).
Using these excerpts, we also generate versions of AI-generated abstracts with respect to their original counterpart, also using GPT-4o. [^9] At the time of conducting this experiment, ChatGPT’s GPT-4o was OpenAI’s premier offering, with traffic surpassing 100 million visits upon its release (PCMag, 2024; Duarte, 2025). Since its inception, ChatGPT has been in use for writing abstracts, generating and editing text, and academic proofreading (Howard et al., 2024; Al Sawi and Alaa, 2024; Brameier et al., 2023). We provide the prompt to generate the abstract and a sample abstract in Figure 15 and Figure 16 (Appendix). Corresponding original abstract from paper is provided in Figure 17 (Appendix).
In this paper, we refer to four versions of abstracts which we use accordingly henceforth. (1) We refer to the authentic human-written abstracts of the published research papers from the peer-reviewed conference publication venues that we used in our study as original abstracts; (2) We refer to the populated AI-generated versions of original abstracts of the respective published papers as AI-generated abstracts; (3) We use the term provided abstracts to indicate those abstracts that we provide the authors to edit. Provided abstracts could be either original abstracts (in Human-noInfo or Human-withInfo treatments) or AI-generated abstracts (in AI-noInfo or AI-withInfo treatments). (4) We use the term edited abstracts to indicate the edited versions of the provided abstracts. To clarify, these edits have been made by the recruited authors. Our investigation includes two participant pools: Authors and Reviewers. We discuss the tasks, incentives, and recruitment procedures of each pool in the respective sections.
3.2. Authors
3.2.1. Task
The main task of the authors is to make edits to the provided abstracts. Authors complete three tasks. In Task 1, we elicit authors’ perceptions about their LLM use. The authors answer four questions. In the first question, we ask if authors use generative AI systems (e.g., ChatGPT) for any writing tasks. Participants choose between four options: they use generative AI to solely generate content, they use generative AI to only edit existing content, they use generative AI to both generate and edit content, or they do not use generative AI for any writing tasks at all. In the second question, we use an incentive-compatible method to elicit authors’ second-order beliefs about the overall use of GPT in writing tasks. We ask what percentage of other authors they believe use GPT to generate new content or edit existing content. Authors are informed that if their selected percentage bracket matches the actual percentage (as calculated by the experimenter), they will receive a bonus of \$0.50. In the third and fourth questions, we ask authors to rate the performance of AI systems in generating and editing content, respectively.
Task 2 entails the main editing task. Every author is presented with three independent research excerpts and their respective provided abstracts. The editing task appears to the authors as illustrated in Figure 9. Figures 3 and 4 provide the detailed illustration provided to the authors to help them understand the instructions. Authors are instructed to first review the excerpts in the left panel of the screen. A scroll bar is provided to the authors to completely review the excerpt, which may extend beyond the scope of the current screen space. Simultaneously, the right panel entails the editing interface that shows the provided abstract. We explicitly instruct the authors to make all edits in the editing interface only. We disabled copy-pasting from the editing interface so that the participants cannot copy and paste to edit on a platform other than the one provided to them. Additionally, this will also limit authors from generating or making edits using existing AI models. The main goal of the authors is to receive at least two Accept decisions by the reviewers on their edited abstracts. Before starting with Task 2, the authors answer two screening questions (Figure 10). These questions were designed to assess the authors’ understanding of task instructions. Correctly answering these questions was a prerequisite for moving forward with the main editing task.
After the screening questions, the authors go through each excerpt-provided abstract combination and make relevant edits. After the authors have made their edits, we elicit two confidence metrics from the authors in an incentive-compatible method. First, we ask the authors to evaluate the provided abstract and indicate how likely they believe it is to receive at least two Accept decisions by the reviewers. In the next question, authors are asked to evaluate the edited abstract, which now includes their own revisions, and indicate how likely they believe it is to receive at least two Accept decisions from reviewers. Responses for both questions are recorded on a slider scale ranging from 0 (not likely at all) to 50 (highly likely). An author edits three instances of provided abstracts and submits respective confidence questions for the three edited abstracts. We also ask authors to rate the quality of the provided abstracts (excellent to poor), describe their frequency of edits, mark the kind of edits authors have made (e.g., addition, deletion, correction, substitution, and reordering), and finally rate the readability of the provided abstracts, using a slider scale where 0 indicates no readability and 100 indicates complete readability.
Authors proceed to Task 3 after they have made their edits and submitted the final versions of edited abstracts. Task 3 involves completing an exit questionnaire that gathers participants’ demographic information (such as gender, education, current profession, age), their first and second-order beliefs about using AI to write abstracts or scientific reports, familiarity AI-generated text versus human-written text, experience with AI tools in daily life, trust in AI tools, and frequency of AI tool usage for writing tasks.
3.2.2. Incentives
Authors earn a fixed participation fee of \$15 for participating in this study and completing all the tasks. Note, authors receive this payment irrespective of their abstracts being accepted by the reviewers or not. Using the methodology of economics experiments, we additionally incentivize the participants to receive a bonus pay of \$15 if a randomly selected abstract out of the three edited abstracts receives at least two Accept decisions from the reviewers. Authors also receive a bonus for their answers in the confidence questions, respectively, according to their submitted confidence levels. E.g., if an author marks 20 on the slider for the edited abstract and it receives at least two Accept decisions from the reviewers, the authors earn a bonus of \$0.20. An author can receive a maximum of \$0.50 from each of the confidence questions.
To encourage genuine effort and careful editing, our incentive structure followed conventions from experimental economics, where performance-contingent rewards are known to elicit higher attention and task quality (Smith, 1982; Camerer and Hogarth, 1999). The fixed participation fee ensured baseline engagement, while performance-based bonuses motivated authors to produce edits that improved clarity and persuasiveness rather than completing the task superficially (Abeler et al., 2011). This helps us maintain data quality and internal validity while aligning with best practices in incentive-aligned experimental and crowd-sourcing studies (Horton and Chilton, 2010; Mason and Watts, 2009). We do not design confidence questions to be an incentive-compatible belief elicitation. Instead, our bonus serves as a small, performance-contingent incentive to increase effort. Since the acceptance is probabilistic, overstating confidence without producing a high-quality edit does not increase expected payoffs.
In total, an author could earn a guaranteed amount of \$15, with the possibility of an additional maximum bonus reward of \$16.50 (i.e., a maximum total of \$31.50). The authors were paid after the completion of the study. After completing all three tasks, the participation fee was immediately transferred to the authors. For the bonus, the computer randomly selected one edited abstract out of the three that the authors submitted for the hypothetical conference set-up, and the authors were informed that they would be paid their bonus based on the reviewer’s decisions for the selected abstract. The average and median time taken for this task were both 1 hour and 8 minutes; our payments justify the stipulated hourly rates on Prolific.
3.2.3. Procedures and Author Recruitment
We recruit 300 authors via the online crowd-sourcing platform Prolific. [^10] This research has been IRB approved by the Office of Responsible Research Practices at Ohio State University under study number 2024E1034. We also pre-registered for our trial at AEA RCT Registry: https://doi.org/10.1257/rct.16740-1.0 We use the Qualtrics interface to set up our study, which was then linked to Prolific for every author. Participation in the study was voluntary. At the onset of the experiment, each author was shown information about the study, followed by an informed consent form.
For our analysis, we exclude 3 authors owing to malformed responses, resulting in 297 authors. Limiting our study to only the CS domain limits us to restricting authors based on their total number of study approvals on Prolific. The median total approval number for our authors is 336.0 approvals, with a minimum of 0 and a maximum of 5560 approvals. 88.51% of authors have their registered first language as English. On average, the authors’ ages (self-reported) range from 25 to 40, and 42.42% of the authors are female (self-reported).
Since all our source material for creating the abstracts is strictly limited to the CS domain, we restrict authors to belonging to the CS domain by prescreening them based on their reported Subjects (proxy for domain of study), as recorded on Prolific. It is a general norm in the highly competitive field of CS that students are engaged in scientific writing since their undergraduate years. To optimize for authors to have some experience with formal writing in CS, we also filter authors by their Highest education level completed. Authors include individuals who have an undergraduate degree (BA/BSc/other), a graduate degree (MA/MSc/MPhil/other), and/or a doctorate (PhD/other). 80.00% of authors are currently maintaining their student status, with 55.56% of authors having at least an undergraduate degree with a major/minor in CS.
Figure 3. An example of the first detailed pictorial representation where experimenters show the authors the author panel and provide them explicit information that edits need to be made directly in the interface. We conducted several pilot studies where we found that the recruited authors were copy-pasting the provided abstract and making edits elsewhere. Providing these graphic instructions and additional screening questions helped mitigate this problem and enabled capturing keystroke-level edits for every abstract.
Figure 4. Illustration of prohibited actions during the abstract editing task. Authors in the second schematic illustration are instructed not to copy the provided abstract into external tools (e.g., text authors, AI assistants), make edits outside the designated experimental interface, or paste modified content back into the system. We inform authors that such actions violate the experimental protocol and may lead to exclusion from the study.
3.3. Editing Interface
Our primary evaluation and analyses are based on recording edits made by the authors. To store these edits, we explored multiple popular text editing tools that provide versioning history, such as Google Docs and Microsoft Word. Having faced several challenges with these tools, we were encouraged to build our own interface. For example, these platforms do not store every keystroke. Instead, they record versions at a fixed time interval, which cannot be modified by the user. This results in us losing many edits that were valuable for our analysis. Furthermore, these platforms do not allow finer-grained control over how the edits can be made. For example, to prevent authors from using AI-assistance to make edits, we needed to disable copying, which none of the existing platforms allow. We also explored Revision History, a Google Docs Extension that offers full document composition record at character-level. This tool offers the magnitude of editing details we sought. However, it operates entirely within Google Docs, is handled online, and therefore does not support exporting the edit history. Google also prohibits automated web scraping of their data, and thus, this pipeline was not reproducible for mass edit log collection.
To build our own interface, we used Firepad, an open-source text editing tool that stores detailed edit history on Firebase Realtime Database and uses CodeMirror as the underlying text editor. While Firepad is no longer being hosted globally, we used its open-source code [^11] https://github.com/FirebaseExtended/firepad to host it ourselves through Firebase Web Hosting. Throughout this study, we used Firepad to record the time, the index, and the edited content per keystroke. We implemented our interface using Firepad 1.5.11, Firebase 5.5.4, and CodeMirror 5.17.0. We plan to open-source this interface for ease of reproduction and use by other researchers. For each author and abstract pair, we create unique Firepad pages. The combination and the index of each pair are used as URL parameters to dynamically populate the authors. Each Firepad page is structured as research findings or excerpts on the left and provided abstracts on the right (Figure 9). To allow authors to refer to the provided abstract when necessary, even though they made changes to the provided abstract, we include the provided abstract at the bottom of the excerpts section. Excerpts, in HTML, are non-editable and embedded in the page. Provided abstracts are pre-loaded in the editing box. We embed the Firepad links into a Qualtrics [^12] https://www.qualtrics.com/ survey form.
3.4. Reviewers
3.4.1. Task
Reviewers complete three tasks. Task 1 for reviewers is identical to the Task 1 completed by the authors. We elicit reviewers’ perceptions about their LLM use. Task 2 entails reviewing abstracts. Each reviewer is presented with a set of 20 abstract pairs—the original abstract that accompanied a selected research paper and the respective edited abstract by an author. [^13] Showing the original abstract in comparison to the edited versions, eliminates any potential bias in reviewers’ abstract acceptance that may arise from the subjective content of the research. We inform the reviewers that the edited abstracts encapsulate the same research idea, with the same research methodology and results as the original abstracts. However, it may or may not be expositionally different from the original abstract. Reviewers need to review the edited abstract in reference to the original abstract and answer the following question: Does the edited abstract provide adequate justice to the research idea presented in the original abstract? The reviewers answer using a slider that ranges from 0 (no justice, worse than the original abstract) to 100 (better than the original abstract), with the default set at 50 (i.e., both abstracts can be considered equally good and one is not better than the other). We also elicit the absolute confidence of the reviewer in each of their 20 decisions. Additionally, to get more insights into the decision-making process of the reviewers, we provide a text box where the reviewers can address this in free-form text. Reviewers proceed to Task 3 after completing Task 2. Task 3 for reviewers, identical to Task 3 completed by the authors, comprises completing an exit questionnaire. We ask two additional questions about the publication record of the reviewers and their prior knowledge of the topics discussed in the abstracts. Note, we do not disclose the source of the edited abstract to the reviewers. [^14] In our setup, we aim to isolate the reviewers’ perception of the quality of edits in the edited abstracts compared to the original. We do not want to confound acceptance decisions with reviewers’ inherent beliefs about AI-generated scientific abstracts.
3.4.2. Incentives
Reviewers earn a fixed participation fee of \$25 for participating in this study and completing all the tasks, irrespective of their decisions in the tasks. Leaning back onto the methodology of incentivized experiments, to increase data quality, we provide an additional bonus of \$8 to the reviewers whose evaluations align closely with those of their peers. One of the twenty abstracts is randomly selected, and two independent reviewers, alongside the reviewer-in-question, assess the same abstract pair using the slider scale described in Task 2. We then compute the median of the three ratings, and if a reviewer’s score falls within 10 points above or below this median, they receive the \$8 bonus. This peer-alignment incentive helps ensure careful, calibrated judgments and reduces random or inattentive responding. Unlike Task 1 incentive for authors, reviewers earn a bonus of \$2 from Task 1, task remaining identical. The first quartile of time taken for the reviewing task is also 1 hour and 8 minutes. However, the reviewers were not required to complete the task during one continuous sitting, and therefore, many took more than a day. Payments to reviewers were made through registered Amazon gift cards.
3.4.3. Procedures and Reviewer Recruitment
We have a total of 159 reviewers. We recruited the reviewers from the current student population at universities who are currently pursuing their graduate and doctoral studies to review 891 abstracts. 174 students signed up for reviewing using the distributed recruitment form. However, only 118 finally took part and completed the study [^15] The first set of 60 reviewers reviewed nine abstracts and received a flat payment of \$15, with a bonus of \$0.50 from Task 1. To maintain a high-quality of the reviewers, we maintained a restricted pool, which required us to increase per-reviewer assignment as well as hourly payment.. 41 remaining reviewers were recruited on Prolific. On Prolific, we maintain the same restrictions for recruiting reviewers as authors; however, the reviewers are now restricted to having completed the highest education level of a graduate degree (MA/MSc/MPhil/other) and/or a doctorate (PhD/other). All reviewers were shown information about the study prior to starting the study, followed by an informed consent form.
32.70% of the reviewers are female (self-reported), 73.59% of reviewers hold at least a master’s degree, 89.94% of reviewers either have a publication or are in progress, with 66.04% having prior knowledge regarding the topics discussed in the abstracts. We have 112 reviewers who provided a Google Scholar or Semantic Scholar link. The median number of publications is 6.5, with a minimum of 1 and a maximum of 61. The median number of citations is 32.5, with a minimum of 0 and a maximum of 4974. For reviewers recruited on Prolific, the median total approval number is 367, with a minimum of 2 and a maximum of 6795 approvals.
Figure 5. Patterns of editing effort in Study 1 (N = 495 abstracts). Left: Distribution of Levenshtein edit distance in two conditions: Human-noInfo (blue) and AI-noInfo (red). Black squares mark the sample means ± 95 % CI; the Welch unequal-variance test indicates a statistically reliable reduction in edits for AI abstracts (p = 0.0293). Middle: Mean edit distance by the authors’ highest education level. Authors with undergraduate or graduate degrees make markedly larger edits to human text than to AI text, whereas doctorate-level editors show the opposite. Right: Scatter-plot of abstract’s perceived readability (0–100) vs the author’s confidence in their edits (0–100). The positive slopes show that authors feel more confident when an abstract reads more smoothly; readability, not AI, as the source explains authors’ higher confidence in the AI-noInfo condition.
3.4.4. Reviewer Assignment
We adopted an automated pipeline to assign abstracts to recruited reviewers by mimicking the in-practice reviewer assignment mechanism in OpenReview, adopted by several CS conferences (Stelmakh et al., 2023). In addition to topical fit, we also try to ensure that no reviewer can review more than one version of an edited abstract originating from the same original abstract to avoid contamination. First, we collect each reviewer’s scholarly profile using the Semantic Scholar APIs. Then we compute the semantic similarity between reviewer publications and base original abstracts using GPT-4o. We account for the reviewer’s citation count to weigh the similarity between an individual paper and submitted abstracts, so that we assign related abstracts closer to the reviewer’s most cited works. Finally, we solve the reviewer assignment as a constrained optimization problem by a minimum-cost flow, with constraints that each submitted abstract gets exactly three reviewers and each reviewer preferably does not review edit versions of the same original abstracts.
3.5. Computed Variables
We additionally compute the following variables for our analysis. For the Author pool, we collect every character-level edit that the author makes to the original abstract provided to them. From here, we compile edit distance score measures the character-level Levenshtein distance [^16] https://en.wikipedia.org/wiki/Levenshtein_distance between the original and edited abstract. This is the main variable of interest. For the reviewers, for each abstract, we collect a score between 0-100 indicating if the edited abstract does justice to the original abstract. If the score is >50>50, then we binarize that decision to be ”Accept” (else ”Reject”) since the original abstract is from a published work. For each edited abstract, three independent ”Accept”/”Reject” decisions are collected, and we compute a final decision (”Accept”/”Reject”) using the majority voting rule. The binarized final decision is the same variable that we use to determine the author bonus.
4. Study A: Human-written vs AI-generated Abstracts
Study A focuses on exploring how authors evaluate and make edits to scientific writing, irrespective of the source information of the script. In a between-subject experiment, we compare two treatments, Human-noInfo and AI-noInfo. Authors are uniformly randomly assigned to either of the two treatment groups. Authors in Human-noInfo treatment edit the original abstracts from the selected published research paper, whereas the authors in the AI-noInfo treatment edit the pre-populated AI-generated abstracts of the selected published research paper. Authors do not have any information about the source of the abstracts and are only provided the excerpts and the abstracts to make their edits. This comparison allows us to observe the natural editing behavior of authors free from disclosure-driven bias. By removing any source cues, we observe how individuals naturally engage with a piece of scientific writing, how they judge its clarity, coherence, and overall quality without being influenced by who (or what) produced it.
4.1. Results
4.1.1. Overall edit
When authors were unaware of whether an abstract was written by a human or an AI system, we found systematic yet heterogeneous differences in editing patterns. We use ordinary least squares (OLS) regression with heteroskedasticity-consistent standard errors to estimate that the authors made significantly smaller edits to AI-generated abstracts compared to human-written ones, on average, about 63 characters fewer (see Figure 5, Left). When the source of the text is unknown, authors perceived AI-written text as closer to a submittable form with lower intent to modify it.
4.1.2. Heterogeneity through education
To account for differences across individual authors, we consider demographic characteristics such as gender and education, along with their interactions with the treatments. Authors with doctoral degrees made substantially more edits to AI-generated abstracts ($p=0.012$; joint Wald test), reversing the average effect. On the contrary, those with undergraduate or master’s degrees made smaller changes (see Figure 5, Middle). This pattern hints that academic (writing) experience promotes the ability to identify idiosyncrasies of AI-generated text. In contrast, less experienced authors interpret the same fluency as high quality, leading to fewer definitive edits.
To probe further, we estimate a mixed-effects model for both author and abstract, capturing unobserved heterogeneity in editing intensity. The model finds negligible between-abstract variance (intra-class correlation =0=0), suggesting that most variation comes from idiosyncratic styles rather than intrinsic differences in abstract quality. In an auxiliary model using perceived readability as a moderator, higher perceived readability strongly predicts lighter editing ($p<0.001$). This does not significantly interact with the AI treatment. We find that the baseline content features of AI-written text shape how authors edit, and that this behavior depends critically on authors’ writing expertise rather than the inherent readability of the content.
Furthermore, authors’ prior familiarity with AI tools, trust in AI, and expectations about AI performance show little influence on how extensively authors edit the abstracts. None of the interaction terms between experience with AI and AI-treatment are statistically significant, suggesting that even frequent AI users or authors who highly trust AI do not edit AI-generated text differently from those less experienced with AI. The main treatment effect remains robust and negative ($\beta=-64.9$, $p=0.038$). Authors make substantially fewer edits to AI-generated abstracts, but the observed differences stem primarily from intrinsic properties of the text (e.g., readability), and not from authors’ prior beliefs on AI.
4.1.3. Ability to distinguish AI-generated text partly explains edits
We next examine whether authors’ self-reported ability to distinguish AI- from human-written text shapes their editing behavior. We focus exclusively on the AI-noInfo condition, since we elicit authors’ ability to distinguish AI vs human-written text in the AI treatment only. We compare the edits of authors who reported being able to distinguish AI from human text (“Yes/Quite often”) with those who reported being unable or unsure (“Rarely/Never”). Authors who claim they could distinguish AI-generated writing made larger edits on average ($M=257.03$, $SD=316.38$, $N=198$) than those who could not ($M=157.56$, $SD=204.57$, $N=27$), the difference being statistically significant ($p=0.033$; Welch’s t-test). To account for potential confounds, we regress edit distance on self-reported ability to distinguish AI text along with original confidence, perceived readability, and demographic controls (gender and education). The effect of AI distinguishability remains positive but becomes statistically insignificant ($\beta=97.15$, $p=0.124$, $N=135$), which indicates the initial difference is partly explained by correlated covariates. Authors with higher pre-edit confidence tend to make smaller edits ($\beta=-6.20$, $p=0.074$), while perceived readability does not significantly predict edit distance ($p=0.990$). This implies that authors who believe they can identify the difference between AI and human writing tend to edit AI-generated text more extensively, but this effect is not independently robust once we control for individual confidence and perceived readability.
4.1.4. Perceived readability mediates confidence
Authors express consistently higher confidence in their edits on AI-generated abstracts ($\beta=2.34$, $p=0.011$ from an OLS) compared to human-generated ones when they remain unaware of the source of the text. However, this effect reduces significantly once authors’ perceived readability is included in the model. When we add perceived readability as a covariate, the coefficient on the AI indicator drops by nearly half and becomes statistically insignificant ($\beta=1.25$, $p=0.123$), while perceived readability itself emerges as a strong positive predictor of confidence ($\beta=0.22$, $p<10^{-18}$), as shown in Figure 5, Right. Authors rate AI-generated abstracts as more readable on average ($M=73.01$, $SD=19.91$) than those written by humans ($M=67.94$, $SD=22.80$), and perceived readability is moderately correlated with confidence ($r=0.48$). Together, these patterns indicate that readability perceptions substantially moderate and explain the higher confidence associated with AI-generated text. Once we account for readability, the direct effect of the AI source largely disappears. Even after including demographic controls (gender and education), perceived readability remains a strong and stable predictor of confidence ($\beta=0.19$, $p<10^{-9}$), whereas the AI effect remains modest and only marginally significant ($\beta=2.24$, $p=0.024$). The perception of how readable the abstract is mediates how the author gains confidence in their edits.
4.1.5. Reviewer decisions
Finally, we model final acceptance decisions for Human-noInfo and AI-noInfo conditions. We find no difference in acceptance probability (AI: $\beta=-0.198$, $p=0.346$), indicating authors are able to bring the abstracts to an acceptable quality when starting from either AI- or human-authored versions when the source is undisclosed.
5. Study B: Source Information Disclosure
The source of text, in recent times, could be human-written or AI-generated. While detection is often challenging, disclosing the source of a scientific text could be fundamental in shaping individual evaluation and interaction patterns with the text. Individuals have often reported reading AI-generated text more critically, anticipating errors, and focusing on the factual accuracy to encounter model hallucinations. In contrast, knowing that a text is human-produced can either invoke trust, or a reluctance to make additional edits. Individuals often feel a stronger sense of oversight when working with AI, or greater deference when editing human work. (Buçinca et al., 2021; Li et al., 2024) provides evidence that transparent source information is crucial to altering calibration, confidence, and task performance.
By varying source disclosure, we can examine how people edit in light of source disclosure and why, whether their actions stem from intrinsic perceptions of quality or from social and cognitive biases linked to authorship cues. We have two disclosure treatments: Human-withInfo and AI-withInfo. In both treatment conditions, we inform authors of the source of the provided abstract. As outlined earlier, the provided abstract could either be the original abstract or the AI-generated abstract.
Figure 6. Treatment effects on editing behavior, heterogeneity, and confidence in Study 2 (N = 495 abstracts). Left: Mean character-level edit distance (± SE) for each treatment condition. Disclosing human authorship (Human-withInfo) significantly reduced edits relative to the no-disclosure baseline (Human-noInfo; $p=.045$), whereas the analogous AI contrast was smaller and nonsignificant. Middle: For AI-generated abstracts, the disclosure effect depends on whether authors say they would personally use generative AI. Among those not willing to use AI, disclosure had little impact; among those willing to use AI, edit intensity fell markedly in the AI-withInfo condition, highlighting that individual AI adoption attitudes shape editing effort. Right: Disclosure of source as AI slightly lowers self-reported pre-edit confidence in acceptance but does not alter the post-edit confidence gains observed across all conditions.
5.1. Results
5.1.1. Overall edit
To examine how disclosure of source information influences the authors’ overall edit intensity, we compare the character-level edit distance pairwise across the final four treatments (Human-noInfo, Human-withInfo, AI-noInfo, and AI-withInfo). Authors make substantially larger edits when they were not told that the abstract was human-written ($M=309.98$, $SD=353.40$, $N=225$). In contrast, than when source information is disclosed ($M=247.87$, $SD=294.86$, $N=219$), there emerges a statistically significant difference, $t(431.96)=2.01$, $p=.045$ (see Figure 6, Left). Disclosure appears to dampen authors’ inclination to edit human-authored text. This proposes a restraint based on self-reputation once authorship is made known. A similar but nonsignificant pattern emerged for AI-generated abstracts. Edits were moderately higher ($p=.176$) in the no-information condition ($M=245.09$, $SD=306.50$, $N=225$) than when AI origin was disclosed ($M=210.50$, $SD=228.27$, $N=222$). As disclosure reduces the total edits, the smaller and nonsignificant effect suggests less sensitivity to provenance cues when participants knew the text was AI-generated. Comparing the two conditions where the source information was disclosed always, authors tend to make larger edits to human-authored original abstracts ($M=247.87$, $SD=294.86$) than to AI-authored ones ($M=210.50$, $SD=228.27$). This difference was, however, not significant, $t(410.46)=1.49$, $p=.138$. Thus, knowing that a text originates from an AI model did not independently increase the extent of editing, once the source information was disclosed. Overall, disclosure mainly affects authors’ effort when the abstract is human-authored. In contrast, knowing that the abstract is AI-generated does not elicit additional editing propensity. We conjecture disclosure primarily acts as a social rather than a purely evaluative signal.
5.1.2. Heterogeneity Effects:
We probe whether the treatment effects on edits vary across author characteristics (gender, age, education, profession) and AI perceptions/usage (e.g., trust in AI, reported use of generative AI, willingness to use AI). For Human-noInfo vs. Human-withInfo (N= [^4] N{=}444), none of the demographic or AI-perception covariates showed a reliable interaction with disclosure conditions. For AI-noInfo vs. AI-withInfo (N=447N{=}447), we observe a significant moderation by author’s willingness to use AI. Specifically, participants who reported they would use AI to edit/generate exhibited a lower editing tendency ($p=0.030$) compared when authors who would not (see Figure 6, Middle). For Human-withInfo vs. AI-withInfo (N=441N{=}441), two AI-perception moderators significantly differentiate the source effect. First, authors who believe others would use GPT show a larger human-versus-AI difference $\beta{=}+193.86$, p=0.005p{=}0.005. Second, authors who reported they would not use GPT exhibit a smaller human-versus-AI difference $\beta{=}{-}189.80$, p=0.028p{=}0.028. This implies that, when the source is always disclosed, these authors amplify their relative editing tendency on human-written abstracts versus AI-generated abstracts. On the other hand, individual reluctance toward AI is associated with a diminished human-vs.-AI gap in editing. Overall, heterogeneity emerges primarily along perceptions about AI, not along demographics. Authors’ own willingness to use AI and their beliefs about others’ AI usage reliably mediate (i) how much AI disclosure affects editing of AI abstracts, and (ii) how strongly authors differentiate between human and AI abstracts when disclosure is present.
5.1.3. Weak treatment effect in Confidence
We examine two measures of authors’ self-reported confidence: (a) original confidence—the expected acceptance of the unedited abstract, and (b) confidence change—the difference between confidence in the edited and original abstract. For each of the three pairwise treatment comparisons, we compute Welch’s unequal-variance t-tests and cluster-robust OLS (clustered by provided abstracts). We then run exploratory mediation analyses for any effect achieving significance $p<.10$. Authors’ confidence does not differ significantly between the Human-noInfo and Human-withInfo conditions. Mean original confidence and post-edit confidence were statistically indistinguishable ($p>.10$), suggesting that disclosing human authorship does not influence how strongly authors believe the abstract would be accepted. For AI-authored abstracts, a modest disclosure effect emerges. Authors show slightly higher original confidence when AI authorship remains undisclosed ($\beta\approx-0.15$, $p=.08$), as well as confidence change did not differ significantly. Thus, disclosure of AI authorship marginally lowers authors’ original confidence in the abstract’s acceptability, though it did not change their post-editing confidence gain. With disclosure (Human-withInfo vs. AI-withInfo), neither original confidence nor confidence change differed significantly ($p>.10$), indicating that once source information is provided, perceived human or AI authorship alone does not affect confidence. Given the marginal effect of AI-disclosure on original confidence, we examine whether perceived readability or beliefs on AI mediate the effect. We find disclosure of AI authorship slightly reduces perceived readability ($p=.06$), and readability positively predicts original confidence ($p=.02$). The Sobel test for the indirect effect remains marginal ($z=1.72,p=.085$). Lower perceived readability may partially mediate the small decline in original confidence when AI authorship is disclosed (see Figure 6, Right). This reduction primarily manifests through perceived readability: when an abstract is identified as AI-written, authors perceive it as slightly less readable, which in turn lowers their confidence. Post-editing confidence gains, however, remain stable across treatments, indicating that authors regain confidence after their edits.
Figure 7. Reviewer decisions differ in AI conditions
5.1.4. Reviewer decisions
We next examine whether the acceptance rate by the reviewers differs across treatment conditions. We model the final decision as a binary variable (1=1=Accept, 0=0=Reject) and compare acceptance rates across the three treatment contrasts, using both $2\times 2$ proportion tests and cluster-robust linear probability models with standard errors clustered by the original abstract. The acceptance rates remain nearly identical, with no statistical significance, across edited abstracts from human treatments: 0.730 in Human-noInfo versus 0.736 in Human-withInfo ($\Delta{=}{-}.006$). For edited abstracts from AI treatments, acceptance rates are 0.706 in AI-noInfo versus 0.773 in AI-withInfo, indicating a modest increase under disclosure ($\Delta{=}{-}.068$, p=0.081p{=}0.081). This suggests that edited versions of AI-generated abstracts, when authors know about the source, make definitive editing decisions that lead to higher acceptance, as shown in Figure 7.
Keeping disclosure effect constant, comparing acceptance of edited abstracts from Human-withInfo (0.736) versus AI-withInfo (0.773) yields no significant difference. We further investigate how AI disclosure to authors may increase acceptance. Focusing on the marginal contrast in acceptance in edited abstracts from (AI-noInfo vs. AI-withInfo), we examine potential mediators reflecting editing effort, confidence, perceived readability, and AI beliefs. For each candidate mediator $M$, we estimate (i) $M{\sim}\text{Treatment}$ and (ii) $\text{FinalDecision}{\sim}\text{Treatment}+M$ using cluster-robust SEs by baseline abstract; the indirect effect was tested via Sobel’s $z$. Among all mediators, editing magnitude through edit distance exhibits the strongest pattern. Disclosure of AI authorship slightly shifts the volume of edits (p<.10p{<}.10), and higher editing effort predicts a higher likelihood of acceptance (p<.05p{<}.05), producing a marginally significant indirect effect. Authors make slightly different editing decisions when aware of AI authorship, and these editorial behaviors modestly enhance the final review outcome.
6. Study C: Quantitative and Qualitative Analyses of Edits
Stylistic guidance for scientific writing is grounded in empirical work on how readers cognitively process text. Psycholinguistics research shows that comprehension improves when information arrives in a predictable fashion. Readability of scientific writing improves when the topic position for context is known to the readers or the stress position for new or important information is expected (Gopen and Swan, 1990; Halliday, 1967). Clarity in scientific prose improves when grammatical subjects mention real “characters” and verbs express actions, rather than nominalizing them (Swales, 2014; Williams and Bizup, 2010). We perform post-hoc quantitative and qualitative analyses of the edits to elicit editing strategies emerging from different treatment conditions.
6.1. Quantitative Stylistic Metrics for Better Scientific Writing
Building on the reader-centric principles of clarity for scientific writing, we operationalize 36 computational linguistic metrics encompassing broad themes of writing: Action, Characters, Cohesion & Coherence, Emphasis, and Sentence Shape. Each metric can be computed with a part-of-speech tagger and first-order dependency labels (e.g., nsubj, ROOT, nsubjpass), making them tractable for post-hoc analysis. We use Spacy [^17] https://spacy.io/ to compute these linguistic features at scale for all provided and edited abstracts to understand how authors edit to improve clarity in their submissions. In subsequent analysis, we find seven of these features stand out that explain the most variance in the data and show significant trends across treatments. Each of these metrics quantify a cognitive signal through abstract’s surface form: 1) Nominalization Density detects when actions are hidden in nouns; 2) Passive-Voice Share reveals disguised agency of the main sentence character; 3) Topic-Continuity Index represents cohesion & coherence by tracking whether successive sentences maintains the same topic; 4) New-at-End Score tests whether novel information is delivered in the stress position; and 5) Subject-Onset Distance, 6) Initial-Momentum, and 7) Pre-Main-Clause Length jointly estimates the sentence shape objective of reaching the subject and verb quickly. All metrics except Topic-Continuity and New-at-End Score trending high are a favorable sign of clear scientific writing.
6.1.1. Stylistic Shifts Across Treatments
We analyze how authors changed the linguistic structure of abstracts across four conditions. For each of 36 stylistic metrics, we computed the change from the original to the edited version ($\Delta=\text{edited}-\text{original}$) and compared conditions using Welch’s unequal-variance $t$-tests. Within each comparison, we controlled for multiple testing.
Across all conditions, there was no reliable stylistic effect of disclosure alone. This indicates that simply knowing whether the text was AI- or Human-generated did not meaningfully change how participants edited its stylistic form.
By contrast, several stylistic differences emerged based on the source of the abstract being edited, as shown in Figure 8. When editing AI-generated versus Human-generated abstracts without disclosure (Human-noInfo vs. AI-noInfo), authors produced lower Topic-Continuity Index ($\Delta_{\text{AI}}-\Delta_{\text{Human}}=+0.051$; $p=0.0002$) and higher Nominalization Density ($\Delta_{\text{AI}}-\Delta_{\text{Human}}=-0.0036$; $p=0.0007$). These effects suggest that authors, when revising human-written texts without source information, tend to make them read heavier and slightly less cohesive, whereas edits to AI-generated abstracts achieve a smoother topical cohesion and fewer nominalizations.
Figure 8. Mean stylistic differences ($\Delta$) between edited and original abstracts across conditions. Shaded bands denote uncertainty distributions around 95% CIs.
Noteworthy stylistic changes appear when authors edited in disclosure conditions (Human-withInfo vs. AI-withInfo). All six metrics show significant changes: Topic-Continuity Index ($p=0.0032$), New-at-End-Score ($p=0.0032$), Passive-Voice Share ($p=0.021$), Initial-Momentum ($p=0.024$), Subject-Onset Distance ($p=0.036$), and Pre-Main-Clause Length ($p=0.044$). Across all six, edits to AI-generated abstracts were to make them structurally shorter, less passive, and more information-dense, while edits to human-written abstracts were to induce longer openings and delayed sentence subjects.
6.2. Edits and AI-detection
To measure the significance of the edits and how AI being the source (perceived or disclosed) moderates editing behavior, we use Pangram [^18] https://www.pangram.com/, an automated AI detection tool, to find cases where author’s edit made the edited abstract more human-written. We can only analyze cases where the provided abstract was AI-generated (i.e., AI-noInfo and AI-withInfo) and use AI detection to label if the edited abstract still can be perceived as “AI-generated.”
We first examine treatment-level differences in the probability that an edited abstract is no longer classified as AI. Authors edited abstracts in the disclosure condition (AI-withInfo) to make the abstract more human-like (23% cases, $N=225$), i.e., can no longer be classified as AI-generated, as compared to no-disclosure condition (AI-noInfo; 18% cases, $N=225$), although this difference is not statistically significant. To characterize this behavior through character-level edits, we find that disclosure significantly increases editing intensity, making edited abstracts more human-like. Conclusively, character-level edit distance emerges as a strong predictor of AI detection outcomes ($p<0.001$).
6.3. Post-experiment Author Interviews
To align our quantitative understanding of how authors edit several treatment conditions, we conducted a series of semi-structured interviews following the main experiment. Studies 1 and 2 quantify how the implicit source and its disclosure moderate editing behavior, whereas these interviews attempt to elicit the reasoning behind such editing behavior. We interviewed 6 authors across all four treatment conditions who successfully completed the editing task. Each interview, conducted over Zoom, lasted for 15 minutes, for which they were compensated with a \$10 flat fee. Post-hoc thematic analysis of the transcripts reveals how authors adopted several editing strategies and their perception of AI-assistance in scientific writing.
Improving readability
Authors across all conditions described a common goal of “making the abstract easier to read.” However, their editing strategies were different depending on treatments. When editing AI‐generated abstracts, authors simplified language and syntax, whereas when editing human‐written abstracts, they reorganized structure and emphasis. One AI–NoInfo editor explained, “I was just thinking how to make it easier for someone else to understand, but still having all the details” (P3). In contrast, authors in Human conditions focused on structure rather than verbosity: “I broke it into paragraphs to make it easier to read” (P1).
Restructuring for emphasis
Beyond clarity, authors re-ordered information to highlight what they saw as the “core” contribution. Those in the Human–withInfo condition displayed the strongest reordering edits, often bringing the paper’s novelty to the beginning. P5 mentioned: “Instead of giving the full background first, I wanted it to say what is being proposed right off the bat,”. AI–WithInfo editors, by contrast, focused on organization: “It would be more legible if restructured into paragraphs that are logically organized” (P4). No-information editors rarely altered structure, limiting their work to grammar or tense adjustments (P0). Disclosure of source thus encouraged rhetorical restructuring rather than surface-level edits.
Perception of authority
Knowing who (or what) authored the abstract reshaped editors’ stance toward authority. Healthy skepticism exists for both types of abstracts; “Even if it’s written by experts, they can also make mistakes” (P5) or “I always have the habit of double-checking because LLMs hallucinate” (P4). Human–WithInfo participants knew experts had written the originals, yet still intervened confidently: “Even if it’s written by experts, they can also make mistakes” (P5). By contrast, authors in no disclosure conditions edited at face value, relying only on textual cues (P0).
Accountability
Across treatments, authors assumed their role to uphold the scientific integrity of the abstracts. They exerted equal oversight to ensure accuracy and coherence, irrespective of human or AI-generated abstracts. As P1 noted, “AI is like an employee—it needs to prove itself before I stop rechecking everything”. Disclosure amplified the sense of accountability: “I still double check anything from AI,” (P4). The editing exercise reveals the author’s intrinsic nature of executing expertise and moral obligation to balance the efficacy and efficiency (by using AI) in the critical act of scientific communication.
7. Discussion
Here we summarize our key findings connecting studies A (Section 4), B (Section 5), and C (Section 6), grounding in prior work and with possible design implications.
Perceived or disclosed source acts as a reputational signal.
Study A shows fewer edits to AI-generated abstracts when the source is undisclosed. Study B shows disclosure (weakly) reverses this pattern, especially reducing edits to human-written abstracts. This result is consistent with prior work showing that source labels activate algorithm aversion and source-identity bias even when content quality is held constant (Dzindolet et al., 2003). Authors explicitly reference deference, accountability, and authority in the interviews. This aligns with evidence in prior work that perceived authorship and authority trigger deference and reduced intervention in evaluative and collaborative settings (Zhu et al., 2025; Fiedler and Döpke, 2025). Collaborative systems that foster AI assistance in scientific writing then must reveal origin of text (e.g., through watermarking) that can assign the appropriate accountability of the authors.
Readability and stylistic edits on AI writing.
Authors express higher perceived readability for AI-generated abstracts in no-disclosure conditions. Still, they make meaningful stylistic edits such as reducing nominalization and improving cohesion. Interview evidence additionally confirms simplification strategies for AI text, also tying back to perceived or disclosed source, acting as a reputational signal. These readability and stylistic patterns reflect broader findings from recent corpus studies showing that AI-generated texts often exhibit distinct linguistic profiles such as more uniform sentence structure and enhanced surface-level coherence (Xu, 2025). AI writing assistants could improve these obvious sloppiness to meaningfully contribute in academic writing.
Expertise mediates trust on AI writing.
Authors with PhD-level experience exhibit completely opposite editing behavior, editing AI-generated abstracts heavily. Mixed-effects models show variance driven by author characteristics rather than abstract quality, which was corroborated by interview narratives reflecting a stronger critical posture among experienced authors. Prior studies similarly show that experts are less susceptible to fluency-driven trust in AI-generated text and are more likely to detect limitations or stylistic artifacts, whereas less experienced users rely more heavily on surface-level coherence as a quality signal (Coutrix and Prost, 2024; Chakrabarty et al., 2024a).
Asymmetric trust on human and AI
Authors report equal responsibility in checking correctness across AI and human text, but more weights put when the text is AI-generated, which was revealed consistently during interviews. Indeed, source disclosure influences how authors edit, hence their trust allocation. When the source is disclosed as AI, we find a strong heterogeneity: authors willing to use AI reduce edits to AI-generated abstracts. On the other hand, with the source being disclosed as human, it substantially reduces the overall edit for human-written abstracts. But, in general, even though editing behavior changes across source information conditions depending on the disclosure, it does not translate into systematic differences in reviewer acceptance. This divergence suggests that disclosure primarily affects authors’ perception of reputation or a manifestation of algorithm aversion, rather than the substantive quality of the final text.
Limitations and future work
Like most experimental studies, our setup also has some limitations which could be addressed in a future work. Though elaborate, our setup is lab-based: a similar study could be done in a real conference setup, where we could observe author behavior with natural incentives. Similarly, capturing editing behavior with real-time AI help, contrary to our offline setup where AI abstracts are already provided. Additionally, confidence and readability plays an important role in describing author’s editing behavior, but we do not incentivize to provide true beliefs to keep the study simple. Finally, This work establishes a foundation for evaluating AI-generated scientific writing in realistic incentivized authorial and review settings. Future research should extend this methodology to other domains beyond computer science to understand domain-specific interactions with AI. Investigating collaborative workflows such as co-writing between multiple authors with AI support would uncover both potential and pitfalls in the use of AI assistance. Additionally, scaling to full-paper evaluations and incorporating multimodal inputs (e.g., tables, figures) would provide a more comprehensive assessment of LLM utility in scientific communication. Finally, data created using such studies could also be used to train models to conduct edits to further improve text quality in the domain of science.
8. Conclusion
We offer a systematic evaluation of AI-assisted scientific writing through a randomized controlled trial simulating a peer review process. Our results show that authorial behavior is significantly influenced by both the implicit source of an abstract and the disclosure of it: authors made the most revisions to human-written texts when source information was withheld, whereas AI-generated abstracts received more extensive edits when attribution was disclosed. However, reviewer decisions were not significantly affected by the abstract’s source. Instead, acceptance outcomes were primarily driven by the degree of meaningful edits. Our analysis suggests that careful edits on the AI-generated abstract can yield higher, if not equal, acceptance as compared to abstracts that are human-authored.
Looking beyond a standalone experiment, longitudinal studies could help us understand how repeated use of AI-assisted writing affects authors’ skill development, stylistic habits, and confidence over time, as well as how it reshapes norms around peer review and scientific communication. Finally, we raise an important point: ethical considerations around transparency, attribution, and responsible use of AI in research are understudied and can have welfare-altering impacts (Hazra et al., 2025). Our study highlights the importance of academic experiences for authors, which play an important role in understanding and effectively utilizing the benefits of AI assistance. At a large scale, we must expand training programs, which are lacking, to inform novice researchers about the socially beneficial use of AI in science. Community-driven investigations of responsible use of AI in science thus must uncover downstream effects on scientific rigor, reproducibility, and community norms to optimally reap the democratic benefits of AI.
GenAI Usage Disclosure
We used GPT-4o to generate AI-generated abstracts central to our experiments. Similarly, GPT-4o was used for assigning reviewers to relevant abstracts for review, again critical for our experiments. We use generative AI-based tools (e.g., Grammarly) to make grammatical checks and grammatical edits for our manuscript.
Acknowledgements.
This material is based upon work supported by the Ai2 Faculty Research Award and the OSU Kirwan Institute. We are grateful to Tuhin Chakraborty, Joseph Chang, Pao Siangliulue, and Peter Clark for feedback on early drafts of this paper. We also thank the anonymous reviewers for their insightful comments.
Appendix A Additional Tables and Figures
Figure 9. This is how the authors view the abstracts. The left non-editable, copy-paste disabled panel includes the excerpts from a selected research paper, and the right panel is the embedded Firepad editor where the authors can make changes.
Figure 10. Screening questions to assess the authors’ understanding of the task.
A.1. Author Comments
We analyze the sentiment of the comments received by the authors. 139 authors out of 297 left a comment, and 54 left a comment indicating “no comment”. Remaining 85 comments are divided into three topics: study (study design), abstract (experiences with the abstracts provided), and AI (opinions regarding the AI usage trend). Table 2 shows the sentiment analysis of each topic. Most reviewers are positive toward the study design, have a less positive opinion toward the abstracts, and are split about AI usage. Word cloud of authors’ comments is in Figure 11.
Table 2. Author comment sentiment analyzed.
Topic Positive Negative Positive Rate
Study 62 9 0.873
Abstract 7 3 0.700
AI 2 2 0.500
Figure 11. Word cloud for authors’ comments.
A.2. Distinguishing AI and Human Texts
We asked the same question to both authors and reviewers: “Can you describe some distinguishing features that help you differentiate between AI-generated and human-written texts?” Popular responses included differences in word choices, sentence structure, and tone of presentation. Some reported that they are barely distinguishable. Word cloud for this question is in Figure 12.
Figure 12. Word cloud for distinguishing factors between AI and human texts, from both authors and reviewers.
Table 3. List of selected papers with venues & years
Title Venue Citation Year
An Analysis of Robustness of Non-Lipschitz Networks NeurIPS 16 2024
RedCode: Risky Code Execution and Generation Benchmark for Code Agents NeurIPS 37 2024
InstructVideo: Instructing Video Diffusion Models with Human Feedback CVPR 62 2024
Locally Adaptive Neural 3D Morphable Models CVPR 3 2024
ToolSword: Unveiling Safety Issues of Large Language Models in Tool Learning Across Three Stages ACL 41 2024
To be Continuous, or to be Discrete, Those are Bits of Questions ACL 4 2024
Prudentia: Findings of an Internet Fairness Watchdog SIGCOMM 5 2024
Rodeo: Making Refinements for Diverse Top-k Queries VLDB 2 2024
Titan: Efficient Multi-target Directed Greybox Fuzzing IEEE S&P 20 2024
C-FRAME: Characterizing and measuring in-the-wild CAPTCHA attacks IEEE S&P 3 2024
An Architecture For Edge Networking Services SIGCOMM 5 2024
ModsNet: Performance-aware Top-k Model Search using Exemplar Datasets VLDB 2 2024
Learning-based Widget Matching for Migrating GUI Test Cases ICSE 29 2024
Harp: Leveraging Quasi-Sequential Characteristics to Accelerate Sequence-to-Graph Mapping of Long Reads ASPLOS 4 2024
GraphTrail: Translating GNN Predictions into Human-Interpretable Logical Rules NeurIPS 7 2024
RainbowCake: Mitigating Cold-starts in Serverless with Layer-wise Container Caching and Sharing ASPLOS 50 2024
Mining Pull Requests to Detect Process Anomalies in Open Source Software Development ICSE 1 2024
Uncertainty-aware Action Decoupling Transformer for Action Anticipation CVPR 17 2024
DocLLM: A Layout-Aware Generative Language Model for Multimodal Document Understanding ACL 91 2024
More is Merrier: Relax the Non-Collusion Assumption in Multi-Server PIR IEEE S&P 8 2024
An exabyte a day: throughput-oriented, large scale, managed data transfers with Effingo SIGCOMM 3 2024
FedSQ: A Secure System for Federated Vector Similarity Queries VLDB 5 2024
Formal Mechanised Semantics of CHERI C: Capabilities, Undefined Behaviour, and Provenance ASPLOS 11 2024
Code Impact Beyond Disciplinary Boundaries: Constructing A Multidisciplinary Dependency Graph and Analyzing Cross-Boundary Impact ICSE 3 2024
Optimistic Critic Reconstruction and Constrained Fine-Tuning for General Offline-to-Online RL NeurIPS 1 2024
DREAM: Diffusion Rectification and Estimation-Adaptive Models CVPR 8 2024
Pareto Optimal Learning for Estimating Large Language Model Errors ACL 5 2024
Distribution Preserving Backdoor Attack in Self-supervised Learning IEEE S&P 37 2024
NegotiaToR: Towards A Simple Yet Effective On-demand Reconfigurable Datacenter Network SIGCOMM 13 2024
Composable Data Management: An Execution Overview VLDB 3 2024
WASP: Workload-Aware Self-Replicating Page-Tables for NUMA Servers ASPLOS 8 2024
Predicting Performance and Accuracy of Mixed-Precision Programs for Precision Tuning ICSE 11 2024
DiffuserLite: Towards Real-time Diffusion Planning NeurIPS 28 2024
DeCoTR: Enhancing Depth Completion with 2D and 3D Attentions CVPR 7 2024
Hard Prompts Made Interpretable: Sparse Entropy Regularization for Prompt Tuning with RL ACL 4 2024
eAUDIT: A Fast, Scalable and Deployable Audit Data Collection System IEEE S&P 31 2024
Fast, Scalable, and Accurate Rate Limiter for RDMA NICs SIGCOMM 6 2024
SpannerLib: Embedding Declarative Information Extraction in an Imperative Workflow VLDB 1 2024
ACES: Accelerating Sparse Matrix Multiplication with Adaptive Execution Flow and Concurrency-Aware Cache Optimizations ASPLOS 4 2024
ACAV: A Framework for Automatic Causality Analysis in Autonomous Vehicle Accident Recordings ICSE 10 2024
ChaosBench: A Multi-Channel, Physics-Based Benchmark for Subseasonal-to-Seasonal Climate Prediction NeurIPS 24 2024
ViLa-MIL: Dual-scale Vision-Language Multiple Instance Learning for Whole Slide Image Classification CVPR 48 2024
Generative Cross-Modal Retrieval: Memorizing Images in Multimodal Language Models for Retrieval and Beyond ACL 37 2024
SoK: The Long Journey of Exploiting and Defending the Legacy of King Harald Bluetooth IEEE S&P 20 2024
OptimusPrime: Unleash Dataplane Programmability through a Transformable Architecture SIGCOMM 3 2024
Figure 13. GPT-4o Prompt we used to extract research findings from papers. The output of the model (after verification) is used as the excerpt to generate the abstract.
Figure 14. An example of research findings fed as input to LLMs to generate abstracts. It is generated using the paper “Locally Adaptive Neural 3D Morphable Models” published at CVPR.
Figure 15. GPT-4o Prompt we used to generate abstract given research findings (excerpts).
Figure 16. An example of AI-generated abstract.
Figure 17. An example of human-written abstract from the published papers.